OpenStudy (anonymous):

7x-3 = 2x -33 Could anyone help me?

OpenStudy (yanasidlinskiy):

7x-3 = 2x-33 -----You need x by itself. You can move either one and will get the same. -7x -7x -3 = -5x - 33 +33 +33 30 = -5x ------Divide by -5 -6 = x. Hope this helps!!:)

OpenStudy (anonymous):

Thank You so much!

OpenStudy (yanasidlinskiy):

Your Welcome!!:) Glad I could help you!!:) Anytime:)